OpenStudy (anonymous):

How do you solve the problem: 2cos^2x=sin2x

OpenStudy (mathmale):

\[2\cos ^{2}x=\sin 2x\]Hint: Find the common identity for sin 2x.

OpenStudy (anonymous):

So you get 2cos^2x = 2sinxcosx then what from there? Because I got that far.

OpenStudy (anonymous):

sin 2x= 2sin x cos x so 2 sin x cos x = 2 cos ² x divide both sides by 2 cos ² x (sin x)/(cos x)=1 tan x=1 x=arctan 1

OpenStudy (mathmale):

How have you (Friesa) previously found angles in situations like this one: tan x = 1 or \[\tan x=\frac{ 1 }{ 1 }?\]

OpenStudy (mathmale):

Note that while it is valid to divide both sides of the equation above by \[2 \cos ² x\] you must also set \[2 \cos ² x = 0\] and solve that for x also.

OpenStudy (mathmale):

Please share your work here. Ask all the questions you want.